An archer shoots arrows at a circular target where the central portion of the target inside is called the bull. The archer hits the bull with probability 1/32. Assume that the archer shoots 96 arrows at the target, and that all shoots are independent. What is an approximated probability that an archer hit not more than one bull?
1
Expert's answer
2020-06-03T19:35:44-0400

This is a binomial distribution with "p=\frac{1}{32},\;n=96."

"P(X\le1)=P(X=0)+P(X=1)=C_{96}^0(\frac{1}{32})^0(\frac{31}{32})^{96}+C_{96}^1(\frac{1}{32})^1(\frac{31}{32})^{95}=0.1944."